OpenStudy (anonymous):

A function is given: f(x) = 3x + 12 a. Determine the inverse of this function and name it g(x). b. Use composite functions to show that these functions are inverses. c. Evaluate f(g(–2)). Explain: What is the domain?

OpenStudy (anonymous):

I need help with a, b, and c please!

OpenStudy (mathstudent55):

a. To find the inverse of a function do this: Write the function as y = 3x + 12 Then swap x and y Then solve for y That function is the inverse of f(x) which you will call g(x)

OpenStudy (anonymous):

(x-12)/3?

OpenStudy (abb0t):

Correct.

OpenStudy (anonymous):

Okay, now how to do b?

OpenStudy (mathstudent55):

A function maps one number of its domain to one number in its range. In your case, for example, f(x) = 3x + 12, if you evaluate the function at x = 1, you get f(1) = 3(1) + 12 = 15, so f(1) = 15. The function maps 1 to 15. If you give the function an input of 1, you get an output of 15. The inverse function takes the value 15 and maps it back to 1. g(x) = (x - 12)/3, so g(15) = (15 -12)/3 = 3/3 = 1. When you evaluate a function at an x value you get a certain y value. If you then evaluate the inverse function at that y value you get back to the original x value. The composition of two functions is to evaluate one function at a certain value, the input, and get the output, and then feed that output as the input of the other function.

OpenStudy (mathstudent55):

Since a function and its inverse do opposite mappings, if you do the composition of a function and its inverse, and evaluate it at any input the final result will be the same as the original input because you applied and "unapplied" the mapping.

OpenStudy (anonymous):

Oh, okay, I got b down now c?

OpenStudy (mathstudent55):

If you can show that fog(x) = x, then you have shown that the functions are inverses.

OpenStudy (anonymous):

fog?

OpenStudy (anonymous):

And so for b, I can just show your examples, and that should cover the answer for that?

OpenStudy (mathstudent55):

Composition of function f and function g. fog(x) = f(g(x)) f(x) = 3x + 12 g(x) = (x - 12)/3